OpenStudy (anonymous):

Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x − 4.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

y = 3x + 7 y = 3x − 7 y = 1 over 3x + 2 y = 1 over 3x − 2

jimthompson5910 (jim_thompson5910):

what is the slope of y = 3x - 4

OpenStudy (anonymous):

i havent done this is in 8 months.. can you walk me through it real quick

jimthompson5910 (jim_thompson5910):

I recommend reading an article like this one http://www.purplemath.com/modules/strtlneq3.htm to refresh your memory

OpenStudy (anonymous):

well i have to leave is 8 mins and its my last question.

OpenStudy (anonymous):

will it take long

jimthompson5910 (jim_thompson5910):

no it shouldn't take long, go ahead and read it

OpenStudy (anonymous):

@jim_thompson5910 is it B?

jimthompson5910 (jim_thompson5910):

what is the slope of y = 3x - 4 ?

OpenStudy (anonymous):

@jim_thompson5910 hey i just got back... is it m=3 b=-4 ?

jimthompson5910 (jim_thompson5910):

m = 3 is the slope the b = -4 is correct, but not needed here

jimthompson5910 (jim_thompson5910):

the parallel line passes through (3,2) so x = 3 and y = 2

jimthompson5910 (jim_thompson5910):

we know: m = 3, x = 3, y = 2 plug them into y = mx+b and solve for b what do you get?

OpenStudy (anonymous):

kk hang on

OpenStudy (anonymous):

so wait by plugging that in i will get b?

jimthompson5910 (jim_thompson5910):

yes and isolating b

OpenStudy (anonymous):

im confused.. @jim_thompson5910

OpenStudy (anonymous):

@sleepyjess

jimthompson5910 (jim_thompson5910):

what do you get when you plugged in the values?

OpenStudy (anonymous):

it said error?

jimthompson5910 (jim_thompson5910):

don't rely on a calculator right now

jimthompson5910 (jim_thompson5910):

simply replace m with 3 replace x with 3 replace y with 2

jimthompson5910 (jim_thompson5910):

you use y = mx+b

OpenStudy (anonymous):

okay so it would look like this... 2=3(3)+B

OpenStudy (anonymous):

?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct, that is equivalent to 2 = 9+b so b = ???

OpenStudy (anonymous):

=-7?

jimthompson5910 (jim_thompson5910):

now we know m = 3 and b = -7 which are then plugged into y = mx+b to get y = 3x - 7

OpenStudy (anonymous):

so i was right!! dang thats so easy but i always make it harder :( thanks man @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome